A car going at speed of 7m/s can be stopped by applying brakes in a shortest distance if 10 m show that the frictional force opposing the motion when brakes are applied is 1/4 th of the weight of the car

Grade:11

2 Answers

Arun
6 years ago
For the motor car;
u = 7 m/s; v = 0; s = 10 m.
Using, v2 = u2 + 2as, we get
 
If we consider, g = 9.8 m/s2, then we get
 
Thus, the resistance to the motion, that is resistive force is
This is one fourth the weight of the car.
SaurabhTrivedi
4 years ago
u = 7m/s , v = 0 ,S=10m v² =u²